A sample from a refuge deposit near the Strait of Magellan has 60% of the carbon 14 of a contemporary living sample. How old was the sample?
Please help! Thank you!!

Answer by Boreal(15235)

P=Poe^rt
P/Po=1/2, which is half as much left
t=5730 years
0.5=e^5730r
ln of both sides
-0.693=5730r
r=-0.0001209
Want P/Po=0.6
0.6=e^(0.0001209)t
ln of both sides
-0.5108= -0.0001209
4224.97 years or 4225 years